# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

### v1

@@ Abstract
The recent success of generative AI highlights the crucial role of high-quality human feedback in building trustworthy AI systems. However, the increasing use of large language models (LLMs) by crowdsourcing workers poses a significant challenge: datasets intended to reflect human input may be compromised by LLM-generated responses. Existing LLM detection approaches often rely on high-dimension training data such as text, making them unsuitable for annotation tasks like multiple-choice labeling. In this work, we investigate the potential of peer prediction — a mechanism that evaluates the information within workers’ responses without using ground truth — to mitigate LLM-assisted cheating in crowdsourcing with a focus on annotation tasks. Our approach quantifies the correlations between worker answers while conditioning on (a subset of) LLM-generated labels available to the requester. Building on prior research, we propose a training-free scoring mechanism with theoretical guarantees under a crowdsourcing model that accounts for LLM collusion. We establish conditions under which our method is effective and empirically demonstrate its robustness in detecting low-effort cheating on real-world crowdsourcing datasets.
@@ 1 Introduction
High-quality human feedback plays a crucial role in shaping modern AI systems, such as ChatGPT, to be trustworthy, safe, and aligned with human preference (Ouyang et al., 2022). This feedback is usually collected through crowdsourcing, where a requester (principal) posts a set of tasks and recruits workers (agents) — who may be trained experts or crowdsourcing workers — to complete them. For example, researchers widely use platforms like Amazon Mechanical Turk and Prolific for tasks such as data labeling and survey administration.
@@
However, the integrity of this data collection process is increasingly threatened by the very AI models it seeks to improve. Evidence suggests that many agents are using large language models (LLMs) to complete tasks such as abstract summarization (Veselovsky et al., 2023) and peer review (Liang et al., 2024), leading to datasets that no longer reflect genuine human input. While LLM-assisted annotations or responses may sometimes be more accurate or coherent than human’s (Gilardi et al., 2023; Törnberg, 2023; Kaikaus et al., 2023), their widespread use clashes with the goal of gathering diverse and authentic human opinions. This issue is particularly concerning in applications that depend on human subjectivity or expertise, such as preference labeling and toxicity detection, where human judgment remains the gold standard.
@@
How to detect potential LLM cheaters and evaluate the quality of a human-contributed dataset? Existing black-box methods for detecting LLM-generated content primarily focus on distinguishing human-written text from AI-generated text by analyzing statistical differences in writing style (Gehrmann et al., 2019; Mao et al., 2024; Koike et al., 2024; Hu et al., 2023). While sometimes effective in open-ended text generation tasks, these training-based approaches are not well-suited for annotation tasks, where the discrete responses lack the rich textual features needed for classification.
@@
In this paper, we develop a theoretically robust scoring metric to evaluate the information in human contributions that is not captured by LLM responses. Our method builds on peer prediction, a scoring mechanism designed to incentivize truthful reporting in the absence of ground truth (Miller et al., 2005; Shnayder et al., 2016). Traditional peer prediction methods evaluate a worker based on the correlation between her responses and those of her peers. Under the standard model, every worker independently observes only high-effort signals — denoted as $X_{i}$ for worker $i$ — after working on a set of i.i.d. questions. Peer prediction mechanisms are shown to be information monotone, meaning that workers maximize their expected score by truthfully reporting their high-effort signals.
@@
However, these methods are ineffective when workers can coordinate on cheap signals, denoted as $Z_{i}$ for worker $i$, that are highly correlated and require minimal effort to produce. In our settings, LLM responses serve as such cheap signals. To address this limitation, we extend peer prediction by measuring correlations between workers’ responses while conditioning on a cheap signal $Z$ that is observable to the requester. For example, $Z_{i}$ and $Z$ could represent responses to the questions independently generated by the same LLM or different LLMs. Intuitively, by conditioning on $Z$, our method scores the information within workers’ responses beyond what can be obtained from LLMs alone, ensuring that human effort is properly rewarded while LLM-assisted cheating is penalized.
@@
Theoretically, we identify the conditions under which our mechanism maintains information monotonicity under two regimes: (1) for any form of cheating strategy and (2) for a subset of practical cheating strategies, which we refer to as naive strategies. Under a naive strategy, rather than reporting honestly on all assigned tasks, the agent sometimes reports noise (random reporting) or copies the LLM output, thereby mixing high-effort signals with uninformative signals. To achieve information monotonicity in the general case, we require that for any two workers $i$ and $j$, $Z_{i}$ (the cheap signal of $i$) is approximately uncorrelated with both $X_{j}$ (the high-effort signal of $j$) and $Z_{j}$ (the cheap signal of $j$), once we condition on $Z$ (see Assumptions 4.1 and 4.2). For naive strategies, a weaker condition suffices: conditioning on $Z$, the correlation between high-effort signals $(X_{i},X_{j})$ is stronger than the correlation between either $(Z_{i},X_{j})$ or $(Z_{i},Z_{j})$ on average (see 4.5).
@@
Empirically, we test our assumptions using two subjective labeling datasets and five types of commonly used LLMs. Our findings show that human responses consistently exhibit stronger correlations than LLM-generated responses when samples of $Z$ and $Z_{i}$ originate from the same model, supporting the validity of 4.5. However, the correlations between independent samples of LLMs responses are usually non-zero, suggesting that Assumptions 4.1 and 4.2 are too strong to hold in practice. This implies that it is feasible to use peer prediction to detect cheating with naive strategies, but detecting arbitrary cheating — where agents choose to cheat even after seeing the high-effort signal — may require substantial ground truth verifications. Finally, under scenarios where 4.5 holds, we simulate cheating agents and evaluate the effectiveness of our method in identifying them. Echoing our theoretical insights, we show that our approach has the most robust performance on mixed crowds compared with all the baselines.
@@ 2 Related Work
Our study is relevant to the usage of LLMs on crowdsourcing platforms. On one hand, if used properly, LLMs can assist better human-AI collaborations or even replace human labor on particular crowdsourcing tasks. For example, Cegin et al. (2023) find that ChatGPT can, to some extent, replace crowdsourcing on paraphrase generation for intent classification tasks. Kaikaus et al. (2023) focus on emotion, sentiment, and cognitive dissonance annotations, where they found that GPT labels are more consistent and reliable than human annotators. Ashktorab et al. (2021) demonstrate that “batch labeling”, a framework of using AI to assist human labeling by allowing a single labeling action to apply to multiple records, can increase the overall labeling accuracy and increase the speed. On the other hand, however, many studies have warned of the dangers of improperly using LLMs in various crowdsourcing tasks. del Rio-Chanona et al. (2023) show that over recent years, LLMs have decreased the activities on Stack Overflow, the leading online Q&A platform for computer programming, by 16%. Regarding the quality of LLM-generated content, evidences exist showing that LLMs tend to have higher bias than human annotators on qualitative analysis (Ashwin et al., 2023), and they tend to exhibit a high degree of unreliability in questions answered in the context of undergraduate computer science. However, LLMs evolve rapidly — a considerable amount of recent work has focused on getting LLMs to emulate human expertise via extensive crowdsourced data (Bach et al., 2022; Zhang et al., 2025; Chiang and Lee, 2023; Köpf et al., 2023). We anticipate that the gap between LLMs and human performance will continue to shrink in the near future.
What’s more, the usage of LLMs in crowdsourcing has uncontrollably increased. In a study, it is reported that as high as 40% of MTurk workers used LLMs to cheat in text summarization tasks (Veselovsky et al., 2023). Furthermore, using a pre-trained LLM, Liang et al. (2024) reported that the use of LLMs in paper writing has significantly increased in recent years — the largest increase in CS papers was as high as 17%. Even worse, an empirical study suggests that humans tend to over-reliant on AI suggestions, even when they are incorrect (Campero et al., 2022). Taking into account both the opportunities and challenges, our work aims to address LLM cheating and foster a healthy and reliable crowdsourcing environment.
Our work builds on advancements in information elicitation, particularly peer prediction. First introduced by Miller et al. (2005), peer prediction mechanisms aim to incentivize truthful reporting without requiring ground truth verification. The core idea is to reward agents based on how well their responses predict those of others. The original mechanism relied on the principal’s knowledge of the correlation between agents’ signals (the information they observe). Later, Dasgupta and Ghosh (2013) extended this idea, removing the need for such knowledge by assuming agents answer multiple i.i.d. binary questions. Subsequent works have expanded this framework to non-binary (but finite) signals (Shnayder et al., 2016; Kong and Schoenebeck, 2019b), infinite signals (Schoenebeck and Yu, 2020), truthful guarantees with fewer questions (Kong, 2020b), heterogeneous agents (Agarwal et al., 2017), adversarial agents (Schoenebeck et al., 2021), and broader cheating strategies (Zhang and Schoenebeck, 2023b).
@@
As a generalization of the correlated agreement (CA) mechanism (Shnayder et al., 2016), our method can score the quality of human responses conditioned on an external information source. We chose the CA mechanism for its simplicity and robust empirical performance, as alternative mechanisms often have complex designs or exhibit high output variance (Burrell and Schoenebeck, 2021; Zhang and Schoenebeck, 2023a; Xu et al., 2024). Our approach is also inspired by Kong and Schoenebeck (2018), who proposed using conditioned mutual information to elicit hierarchical information. Unlike their framework, which assumes an information hierarchy where more effort yields strictly more informative signals, our method does not require hierarchical structures or sequential elicitation. Instead, we assume the principal has access to samples of low-effort information and focus on empirically verifying the conditions under which our method works (or not). Lastly, we propose a heuristic generalization of our method to handle textual and other forms of complex data (Appendix C). This approach is related to the discussion of applying peer prediction to elicit textual information using LLMs (Lu et al., 2024; Wu and Hartline, 2024).
@@
The studies on LLM-content detection are also related to this work. Common approaches for LLM content detection include watermarking (Gu et al., 2022; Kirchenbauer et al., 2023; Lee et al., 2023; Zhao et al., 2023; Yang et al., 2023a), which embeds identifiable markers into data, models, or generated text; statistical detection, which leverage differences in output by analyzing logits statistics (Su et al., 2023; Vasilatos et al., 2023) or vocabulary statistics (Gehrmann et al., 2019; Mao et al., 2024; Yang et al., 2023b); and deep learning-based methods, which train models to distinguish AI-generated content from human-written text (Koike et al., 2024; Hu et al., 2023; Shah et al., 2023). For a more comprehensive review, we refer to the survey paper (Wu et al., 2024). However, these methods typically rely on abundant textual data for training and detection. In contrast, our approach is training-free and addresses cheating in scenarios where the critical distinction between human-written and AI-generated content lies in the information level rather than the vocabulary level. For instance, our method is designed for tasks like answering multiple-choice questions, where no direct textual response is available.
@@ 3 Model
Consider a truth-discovery setting where there are $m$ i.i.d. questions. The principal distributes these questions to $n$ agents such that each agent answers multiple questions and each question is answered by more than one agent. We first present the information structure when two agents, denoted as $i$ and $j$, answer an arbitrary question. Because questions are i.i.d., the same information structure will be applied to any question.
@@
Signals  Suppose the underlying ground truth of the question is $Y\in\Sigma$. For example, if the answer to the question is either “yes” or “no”, then $|\Sigma|=2$. Suppose agent $i$ and $j$ both answer this question. After careful thinking, they each observe a signal $X_{i}$ and $X_{j}$ respectively. For simplicity, we assume the signal space is the same as the ground truth space, i.e. $X_{i},X_{j}\in\Sigma$. We can interpret $X_{i}$ and $X_{j}$ as the high-effort signals, which are agents’ best guess of $Y$ after exerting a substantial amount of effort.
In addition to their high-effort signals, agents can obtain cheap signals $Z_{i},Z_{j}\in\Sigma$ that may or may not depend on $Y$. The cheap signals are usually strongly correlated, sometimes even more correlated than agents’ high-effort signals. For example, these can be the responses of two (potentially different) LLMs.
We present the causal relationship of these variables in Fig. 1. Note that $X_{i}$ and $Z_{i}$ are independent conditioned on $Y$, meaning that the high-effort signal is only correlated with the cheap signal via the ground truth. However, $Z_{i}$ and $Z_{j}$ may have additional correlations conditioned on $Y$. This may happen when agents cheat with LLMs trained on similar data.
Strategies  The principal’s goal is to collect high-effort signals from agents. This means that we consider scenarios where human expertise remains the gold standard, or AI underperforms human agents. Such situations arise when the notion of ground truth is inherently subjective, shaped by human preferences — such as when labelers are asked to rank LLM responses according to personal preference, a common practice in reinforcement learning with human feedback.
However, agents can strategically report based on their high-effort signal $X_{i}$ and the cheap signal $Z_{i}$. Mathematically, the report $\hat{X}_{i}$ is a (random) function of $X_{i}$ and $Z_{i}$, denoted as $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ where $\theta_{i}\in\Theta$. Let $\tau$ be the truth-telling strategy, i.e. $\tau(X_{i},Z_{i})=X_{i}$, and let $\mu_{i}$ denote an uninformative strategy such that $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$. The definitions for $\hat{X}_{j}$ and $\theta_{j}$ are analogous.
We highlight a special type of strategy, called the naive strategy, which is more natural and common on crowdsourcing platforms.
@@
A naive strategy $\nu\in\Theta_{\nu}$ is a convex combination between truth-telling and an uninformative strategy $\mu$, i.e.
@@
Examples of naive strategies include: exerting effort and reporting $X_{i}$ for some questions while randomly selecting answers for the rest; or fully relying on an LLM and always reporting $Z_{i}$. In contrast, non-naive strategies are typically complex and impractical, often resembling malicious behavior rather than mere low-effort cheating. For example, an agent may first exert effort to obtain $X_{i}$ for each question and then mix it with $Z_{i}$ by reporting “yes” if either $X_{i}$ or $Z_{i}$ is “yes”, and “no” otherwise.
Principal Information  The principal can observe the reports $\hat{X}_{i}$ and $\hat{X}_{j}$, but not the underlying signals $X_{i},X_{j}$ and $Z_{i},Z_{j}$. Furthermore, the principal cannot directly evaluate human responses using the ground truth $Y$. Without knowing $Y$, the principal cannot distinguish the real information $X_{i}$ and the cheap signal $Z_{i}$ without further assumptions. Our method assumes that the principal can observe a signal $Z$ that is (strongly) correlated with the agents’ cheap signals. In the context of LLM cheating, for example, the principal can generate answers to a fraction of the questions using a popular LLM.
The effectiveness of our method necessarily depends on the quality of $Z$ — how well the principal’s information can block the correlation between agents’ cheap signals. We will introduce our method and the assumptions in Section 4 and empirically verify the assumptions in Section 5.3.

### v2

# Evaluating LLM-Contaminated Crowdsourcing Data Without Ground Truth

@@ Abstract
The recent success of generative AI highlights the crucial role of high-quality human feedback in building trustworthy AI systems. However, the increasing use of large language models (LLMs) by crowdsourcing workers poses a significant challenge: datasets intended to reflect human input may be compromised by LLM-generated responses. Existing LLM detection approaches often rely on high-dimensional training data such as text, making them unsuitable for structured annotation tasks like multiple-choice labeling. In this work, we investigate the potential of peer prediction—a mechanism that evaluates the information within workers’ responses—to mitigate LLM-assisted cheating in crowdsourcing with a focus on annotation tasks. Our approach quantifies the correlations between worker answers while conditioning on (a subset of) LLM-generated labels available to the requester. Building on prior research, we propose a training-free scoring mechanism with theoretical guarantees under a novel model that accounts for LLM collusion. We establish conditions under which our method is effective and empirically demonstrate its robustness in detecting low-effort cheating on real-world crowdsourcing datasets.
@@ 1 Introduction
High-quality human feedback plays a crucial role in shaping modern AI systems, such as ChatGPT, to be trustworthy, safe, and aligned with human preference ouyang2022training. This feedback is usually collected through crowdsourcing, where a requester posts tasks and recruits trained experts or general participants to complete them. For example, platforms like Amazon Mechanical Turk and Prolific are widely used by researchers for tasks such as data labeling and survey administration.
@@
However, the integrity of this data collection process is increasingly threatened by the very AI models it seeks to improve. Evidence suggests that many agents rely on large language models (LLMs) to complete tasks such as abstract summarization (veselovsky2023artificial) and peer review (liang2024mapping), leading to datasets that no longer reflect genuine human input, a phenomenon known as LLM contamination. While LLM-assisted annotations or responses may sometimes be more accurate or coherent than human’s (Gilardi_2023; törnberg2023chatgpt4; 10386425), their widespread use undermines the fundamental goal of gathering diverse and authentic human opinions. This issue is particularly concerning in applications that depend on human subjectivity or expertise, such as preference labeling and toxicity detection, where human judgment remains the gold standard.
@@
How to detect potential LLM-misuse and evaluate the quality of a human-contributed dataset? Existing black-box methods for detecting LLM-generated content primarily focus on distinguishing human-written text from AI-generated text by analyzing statistical differences in writing style (gehrmann2019gltr; mao2024raidar; koike2024outfox; hu2023radar). While effective in open-ended text generation tasks, these training-based approaches are not well-suited for structured annotation tasks like multiple-choice labeling, where the discrete responses lack the rich textual features needed for classification.
@@
In this paper, we develop a theoretically robust scoring metric to evaluate the quality of human contributions in addition to existing LLM answers, effectively distinguishing workers with less informative responses. Our method builds on peer prediction, a scoring mechanism designed to incentivize truthful reporting in the absence of ground truth miller2005eliciting; shnayder2016informed. Traditional peer prediction methods evaluate a worker based on the correlation between her responses and those of her peers. In the standard model, every worker independently observes only high-effort signals, denoted as $X_{i}$ for worker $i$, after working on a set of i.i.d. questions. Peer prediction mechanisms are shown to be information monotone, meaning that workers maximize their expected score by truthfully reporting their high-effort signals.
@@
However, these methods are ineffective when workers can coordinate on cheap signals, denoted as $Z_{i}$ for worker $i$, that are highly correlated and require minimal effort to produce. In our settings, LLM responses serve as such cheap signals. To address this limitation, we extend peer prediction by measuring correlations between workers’ responses while conditioning on a signal $Z$ that is observable to the requester. For example, $Z_{i}$ and $Z$ could represent responses to the questions independently generated by the same LLM or different LLMs. Intuitively, by conditioning on $Z$, our method scores the information within workers’ responses beyond what can be obtained from LLMs alone, ensuring that human effort is properly rewarded while AI-reliant is penalized.
@@
Theoretically, we identify the conditions under which our mechanism maintains information monotonicity under two regimes: (1) for any form of manipulation strategy and (2) for a subset of practical cheating strategies, which we refer to as lazy-reporting strategies. A lazy-reporting strategy, for instance, includes blindly relying on LLM on a fraction of the tasks while answering the rest truthfully. To achieve information monotonicity in the general case, for an arbitrary pair of workers, $i$’s cheap signal $Z_{i}$ has to be approximately uncorrelated with $j$’s high-effort signal $X_{j}$ and their cheap signal $Z_{j}$, conditioned on $Z$ (Assumptions 4.1 and 4.2). For lazy-reporting strategies, a weaker condition suffices: conditioning on $Z$, the correlation between high-effort signals $(X_{i},X_{j})$ is stronger than the correlation between both $(Z_{i},X_{j})$ and $(Z_{i},Z_{j})$ (4.5).
@@
Empirically, we test our assumptions using two subjective labeling datasets and five types of commonly used LLMs. Our findings show that human responses consistently exhibit stronger correlations than LLM-generated responses when samples of $Z$ and $Z_{i}$ originate from the same model, supporting the validity of 4.5. However, the correlations between independent samples of LLMs responses are usually non-zero, suggesting that Assumptions 4.1 and 4.2 are too strong to hold in practice. Finally, under scenarios where 4.5 holds, we evaluate the effectiveness of our method in detecting low-effort agents. Echoing our theoretical insights, we show that our approach has the most robust performance on mixed crowd compared with all the baselines.
@@ 2 Related Work
Our study relates to the studies on LLMs usage in crowdsourcing. cegin2023chatgpt find that ChatGPT can partially replace human workers in paraphrase generation, and 10386425 report that GPT labels are more consistent than humans in sentiment annotations. ashktorab2021ai demonstrate that “batch labeling”, a framework of using AI to assist human labeling by allowing a single labeling action to apply to multiple records, can increase the overall labeling accuracy and increase the speed. However, several studies also highlight potential downsides: del2023large report a 16% decline in Stack Overflow activity following the rise of LLMs, and ashwin2023using find that LLMs exhibit greater bias than human annotators in qualitative analysis. In this context, our paper introduces a scoring mechanism that quantitatively measures the informativeness of human reports relative to the information already provided by AI.
@@
Our work builds on advancements in information elicitation, particularly peer prediction. As a generalization of the correlated agreement (CA) mechanism (shnayder2016informed), our method can score the quality of human responses conditioned on an external information source. We chose the CA mechanism for its simplicity and robust empirical performance (burrell2021measurement; zhang2022high). Our approach is also inspired by kong2018eliciting, who proposed using conditioned mutual information to elicit hierarchical information. Their framework assumes an information hierarchy where more effort yields strictly more informative signals, so low-effort signals can be elicited first, which are used to further elicit high-effort signals. In contrast, our method assumes that the principal has access to samples of low-effort signals, which is more straightforward and practical in the LLM-influenced crowdsourcing settings. Lastly, we propose a heuristic generalization of our method to handle textual data (Appendix F), which is related to methods that elicit textual information using LLMs (textualpp2024; wu2024elicitationgpt).
@@
The studies on LLM-content detection are also related. Common approaches for LLM content detection include watermarking (gu2022watermarking; lee2023wrote; zhao2023provable; yang2023watermarking), which embeds identifiable markers into data, models, or generated text; statistical detection, which leverage differences in output by analyzing logits statistics (su2023detectllm; vasilatos2023howkgpt) or vocabulary statistics (gehrmann2019gltr; mao2024raidar); and deep learning-based methods, which train models to distinguish AI-generated content from human-written text (koike2024outfox; hu2023radar; shah2023detecting). However, these methods typically rely on abundant textual data for training and detection. In contrast, our approach is training-free and addresses LLM-contamination in non-textual responses scenarios, such as answering multi-choice questions.
@@ 3 Model
Consider a truth-discovery setting where there are $m$ i.i.d. questions. The principal (requester) distributes these questions to $n$ agents (workers) such that each agent answers multiple questions and each question is answered by more than one agent. We first present the information structure when two agents, denoted as $i$ and $j$, answer an arbitrary question. Because questions are i.i.d., the same information structure will be applied to any question.
@@
Signals  Suppose the underlying ground truth of the question is $Y\in\Sigma$. For example, if the answer to the question is either “yes” or “no”, then $|\Sigma|=2$. Suppose agents $i$ and $j$ both answer this question. After exerting effort, they each observe a signal $X_{i}$ and $X_{j}$ respectively. For simplicity, we assume the signal space is the same as the ground truth space, i.e. $X_{i},X_{j}\in\Sigma$. We can interpret $X_{i}$ and $X_{j}$ as the full-effort signals that the agents can obtain which are their best guesses of $Y$.
In addition to full-effort signals, agents can obtain cheap signals $Z_{i},Z_{j}\in\Sigma$ that may or may not depend on $Y$. Cheap signals are usually mutually correlated with each other, sometimes even more correlated than agents’ full-effort signals. For example, these can be the responses of two (potentially different) LLMs.
We present the causal relationship among these variables in Figure 1. Note that we assume $X_{i}$ and $Z_{i}$ are independent conditioned on $Y$, meaning that the full-effort signal is only correlated with the cheap signal via the ground truth. However, $Z_{i}$ and $Z_{j}$ may have additional correlations conditioned on $Y$. This may happen when agents rely on LLMs trained on similar data.
Strategies  The principal’s goal is to recover the ground truth as accurately as possible. We consider scenarios where AI underperforms human agents, motivating the principal to seek full-effort information from human agents. Such situations arise when the notion of ground truth is inherently subjective, shaped by human preferences—such as when labelers are asked to rank LLM responses according to personal preference, a common practice in reinforcement learning with human feedback.
However, agents can strategically report based on their high-effort signal $X_{i}$ and the cheap signal $Z_{i}$. Mathematically, the report $\hat{X}_{i}$ is a (random) function of $X_{i}$ and $Z_{i}$, denoted as $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ where $\theta_{i}\in\Theta$. Let $\tau$ be the truth-telling strategy, i.e. $\tau(X_{i},Z_{i})=X_{i}$, and let $\mu_{i}$ denote a no-effort strategy such that $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$ conditioned on $Z_{i}$. The argument for $\hat{X}_{j}$ and $\theta_{j}$ are analogous.
We highlight a special type of strategy, called the lazy-reporting strategy, which is more practical and common on crowdsourcing platforms.
@@
A lazy-reporting strategy $\nu\in\Theta_{\nu}$ is a convex combination between truth-telling and a no-effort strategy $\mu$, i.e.
@@
In particular, $\nu$ does not make $\hat{X}_{i}$ depend jointly on both $X_{i}$ and $Z_{i}$ on the same question. Examples of lazy-reporting strategies include: exerting effort and reporting $X_{i}$ for some questions while randomly selecting answers for the rest; or fully relying on an LLM and always reporting $Z_{i}$. In contrast, non-lazy reporting strategies are typically complex and impractical, often resembling malicious rather than mere low-effort behavior. For example, an agent may first exert effort to obtain $X_{i}$ for each question and then mix it with $Z_{i}$ by reporting “yes” if either $X_{i}$ or $Z_{i}$ is “yes”, and “no” otherwise.
Principal Information  The principal can observe the reports $\hat{X}_{i}$ and $\hat{X}_{j}$, but not the underlying signals $X_{i},X_{j}$ and $Z_{i},Z_{j}$. This implies that the principal cannot directly evaluate human responses using the ground truth $Y$. Without knowing $Y$, the principal cannot distinguish the real information $X_{i}$ and the cheap signal $Z_{i}$ without further assumptions. Our method assumes that the principal can observe a signal $Z$ that is (strongly) correlated with the agents’ cheap signals. In the context of LLM contamination, for example, the principal can generate answers to a fraction of the questions using a popular LLM.
The effectiveness of our method necessarily depends on the quality of ZZ—how well the principal’s information can block the correlation between agents’ cheap signals. We will introduce our method and the assumptions in Section 4 and empirically verify the assumptions in Section 5.2.